# Tailored Teaching with Balanced Difficulty: Elevating Reasoning in Multimodal Chain-of-Thought via Prompt Curriculum

### v1

@@ Introduction
Baselines and Model Variants.
In our experiment, the following five methods are used as the main baselines: Chain-of-thought (ZS-CoT) (Wei et al. 2022), Few shot CoT (FS-CoT), Self-consistency (Self-Con) (Wang et al. 2022), Auto-CoT (Zhang et al. 2022) and Active Prompt (Active-Pro) (Diao et al. 2023). FS-CoT uses the same annotation process as our method, the only difference being that it randomly selects questions from the training data for annotation. We select multiple cutting-edge MLLMs as experimental models, including llama3.2-vision:11b, llava:7b, and Qwen2.5-VL:7b.
Implementation Details
Hyperparameters.
For each dataset, we set the number of prompt examples to 4. In the training set sampling phase, we set the number of sampling iterations to 5. During both the sampling and the inference phases of the experiments, the temperature is uniformly set at $0.5$. Unless otherwise specified, the multimodal large language models used in the experiments are llama3.2-vision:11b, llava:7b, and Qwen2.5-VL:7b.
Uncertainty Assessment.
In the experimental process, we adopt a zero-shot sampling strategy, which does not rely on additional examples or guidance information. For ScienceQA, A-OKVQA, and VQA, we perform sampling on the complete training sample sets, while for the VQAv2 and TextVQA datasets, we only sample 10,000 samples. For the sampling frequency parameter, we set $K=5$. We consistently use “Disagreement” [^1] In preliminary experiments, variance and entropy exhibit suboptimal performance in multimodal scenarios.—a more intuitive and accurate method—as the uncertainty metric.
Constructing Examples.
We focus on the innovating and optimizating strategies for prompt example selection. To reduce human labor as much as possible, we eliminate the need for manual annotation of prompt examples. Instead, we construct prompt instances by directly concatenating the question, reasoning, and answer from the dataset. Details are provided in Appendix A.
@@
Main Results
For Q1: CAMS consistently outperforms nearly all baseline methods.
For Q2: CAMS can greatly reduce accuracy instability caused by randomly selecting prompt examples.
METHOD Datasets Avg.
ScienceQA A-OKVQA OK-VQA VQAv2 TextVQA
Llama3.2-vision:11b
ZS-CoT 38.08 47.42 28.74 57.45 37.52 41.84
Unc-Eva 63.40 56.68 43.65 59.75 61.32 56.96
Com-Eva 61.92 55.92 48.10 60.66 60.40 57.40
Ours 68.22 59.39 51.89 61.89 62.76 60.83
For Q3: Both the uncertainty analysis and complexity evaluation modules can effectively improve accuracy.
For Q4: CAMS can improve the model’s accuracy in subdivided domains.
To further investigate the efficacy of CAMS in specific subfields, we conduct systematic experiments on the ScienceQA dataset. Specifically, based on disciplinary attributes, the ScienceQA dataset is categorized into three major groups: natural sciences (NAT), social sciences (SOC), and linguistic sciences (LAN). According to difficulty gradients, it is also divided into two levels: grades 1–6 (G1-6) and grades 7–12 (G7-12), thus constructing multidimensional subscenarios that capture both domain specificity and cognitive complexity. As shown in Figure 6, in the five subfields mentioned above, CAMS almost universally outperforms the three baseline methods - Self-con, ZS-CoT, and FS-CoT - in terms of precision, especially in the NAT, SOC, and G1-6 domains. Whether in knowledge-centric scenarios emphasizing disciplinary dimensions or in learning settings focused on difficulty levels, CAMS demonstrates significant advantages, confirming its ability to substantially improve model accuracy within specialized domains.
For Q5: A selection strategy that combines easy and difficult examples is more beneficial to the model.
@@
Conclusion and Future Work

### v2

@@ 4.3 Main Results
Module Datasets Avg.
Unc-Eva Com-Eva ScienceQA A-OKVQA OK-VQA VQAv2 TextVQA
Llama3.2-vision:11b
(a) ZS-CoT 38.08 $\downarrow 48.58\%$ 47.42 $\downarrow 20.15\%$ 28.74 $\downarrow 44.61\%$ 57.45 $\downarrow 7.17\%$ 37.52 $\downarrow 40.22\%$ 41.84 $\downarrow 31.22\%$
(b) 63.40 $\downarrow 7.07\%$ 56.68 $\downarrow 4.56\%$ 43.65 $\downarrow 15.88\%$ 59.75 $\downarrow 3.46\%$ 61.32 $\downarrow 2.29\%$ 56.96 $\downarrow 6.36\%$
(c) 61.92 $\downarrow 9.23\%$ 55.92 $\downarrow 5.84\%$ 48.10 $\downarrow 7.30\%$ 60.66 $\downarrow 1.99\%$ 60.40 $\downarrow 3.76\%$ 57.40 $\downarrow 5.64\%$
(d) 68.22 59.39 51.89 61.89 62.76 60.83
4.3.3 For Q3: Both the uncertainty analysis and complexity evaluation modules can effectively improve accuracy.
4.3.4 For Q4: CAMS can improve the model’s accuracy in subdivided domains.
To further investigate the efficacy of CAMS in specific subfields, we conduct systematic experiments on the ScienceQA dataset. Specifically, based on disciplinary attributes, the ScienceQA dataset is categorized into three major groups: natural sciences (NAT), social sciences (SOC), and linguistic sciences (LAN). According to difficulty gradients, it is also divided into two levels: grades 1–6 (G1-6) and grades 7–12 (G7-12), thus constructing multidimensional subscenarios that capture both domain specificity and cognitive complexity. As shown in Figure 4, in the five subfields mentioned above, CAMS almost universally outperforms the three baseline methods - Self-con, ZS-CoT, and FS-CoT - in terms of precision, especially in the NAT, SOC, and G1-6 domains. Whether in knowledge-centric scenarios emphasizing disciplinary dimensions or in learning settings focused on difficulty levels, CAMS demonstrates significant advantages, confirming its ability to substantially improve model accuracy within specialized domains. Meanwhile, this also demonstrates the generalization capability of the CAMS framework.
@@
4.3.5 For Q5: A selection strategy that combines easy and difficult examples is more beneficial to the model.
@@
5 Conclusion and Future Work
@@
Reproducibility statement
We have submitted the relevant code in the supplementary materials. The names of the experimental benchmarks, the prompt templates used, and the model’s hyperparameter settings can all be found in The Appendix A.1,  A.2,  A.3 and Table 3. Section 4 (Experiments) provides a detailed description of the experimental setup for the mechanism experiments.
Appendix A Appendix
A.1 The construction of Prompt
We construct prompt examples by integrating questions, reasoning steps, and answers from the dataset:
$E=\{(q_{1},c_{1},a_{1}),(q_{2},c_{2},a_{2}),...,(q_{n},c_{n},a_{n})\}$ (7)
where $q$ denotes questions, $c$ denotes reasoning steps, $a$ denotes answers, and $n$ denotes the number of examples.
A.2 Comparison between CAMS and Auto-CoT
Figure 6 demonstrates that the prompt examples selected by CAMS can, to a certain extent, enhance the strengths and mitigate the weaknesses of multimodal large models, thereby better facilitating their completion of reasoning tasks.
Figure 6: Comparative cases between CAMS and Auto-CoT
A.3 Prompt Demonstration
Details are provided in Table 3. It is important to note that the captions in the examples are replaced with actual images in real-world applications; text is used here solely for the convenience of demonstration.
A.4 The Use of Large Language Models (LLM)
In order to enhance the language quality and clarity of this academic paper, the author utilized AI-powered tools for text refinement during the writing process. The specific details are as follows:
Purpose of Use: The primary purposes for using AI tools were to:
  • Check grammar and spelling for certain sentences.
  • Optimize vocabulary choices for more precise and academic expression.
  • Adjust sentence structures to improve logical coherence and readability between paragraphs.
Method of Use: The author input original paragraphs written by themselves into the AI tools and then manually judged, filtered, and revised the text based on the refinement suggestions provided. All adopted changes were carefully considered by the author to ensure they fully align with the original intent and academic rigor of the paper.
Disclaimer of Responsibility: All academic content in this paper, including core arguments, research data, result analysis, argumentation process, and final conclusions, was independently created and is the sole responsibility of the author. The AI tools were used purely as an auxiliary aid and did not generate any critical academic viewpoints, research data, or conclusions. The author assumes full responsibility for the final content of the paper.
Tools Used: The AI tool used in this process is: GPT-4.
Table 3: Prompt template of query formulation.
SYSTEM:
You are a professional VQA task solver
USER:
Given the problem and its related images, you need to generate answers for the VQA task.
The generated answers must be derived from direct visual observation of the image—do not include speculative content, assumptions, or information not visible in the image.
When answering questions: if the image clearly shows the required information, output a specific, concise answer; if the image does not contain enough information to answer (e.g., object not visible, detail unclear), output “insufficient information”; if the image contradicts the question (e.g., question asks about a “red car” but image shows a blue car), output a negative answer with correct visual details.
Do not answer subjective judgment questions (e.g., “Is the image interesting?”)—visual facts only.
You MUST only respond in the format as described below.
DO NOT RESPOND WITH ANYTHING ELSE.
Response Format:The answer is …, because … .
Here are an example:
“question”:“Which of these organisms contains matter that was once part of the phytoplankton?”
“choices”:[“black rockfish”, “sea otter”]
“hint”:“Below is a food web from an ocean ecosystem in Monterey Bay, off the coast of California. A food web models how the matter eaten by organisms moves through an ecosystem. The arrows in a food web represent how matter moves between organisms in an ecosystem.”
“caption”:“A painting of a penguin on a wall.”
“lecture”:“A food web is a model. A food web shows where organisms in an ecosystem get their food. Models can make things in nature easier to understand because models can represent complex things in a simpler way. If a food web showed every organism in an ecosystem, the food web would be hard to understand. So, each food web shows how some organisms in an ecosystem can get their food. Arrows show how matter moves. A food web has arrows that point from one organism to another. Each arrow shows the direction that matter moves when one organism eats another organism. An arrow starts from the organism that is eaten. The arrow points to the organism that is doing the eating. An organism in a food web can have more than one arrow pointing from it. This shows that the organism is eaten by more than one other organism in the food web. An organism in a food web can also have more than one arrow pointing to it. This shows that the organism eats more than one other organism in the food web.”
“output”:“The answer is A, because Use the arrows to follow how matter moves through this food web. For each answer choice, try to find a path of arrows that starts from the phytoplankton. The only arrow pointing to the sea otter starts from the sea urchin. The only arrow pointing to the sea urchin starts from the kelp. No arrow points to the kelp. So, in this food web, matter does not move from the phytoplankton to the sea otter.There are two paths matter can take from the phytoplankton to the plainfin midshipman: phytoplankton>->plainfin midshipman. phytoplankton>->zooplankton>->plainfin midshipman. There is one path matter can take from the phytoplankton to the black rockfish: phytoplankton>->zooplankton>->black rockfish. There is one path matter can take from the phytoplankton to the zooplankton: phytoplankton>->zooplankton.”
Now complete your output with following the above rules.
Input:
Output:
Algorithm 1 CAMS
0: Training set $D=\{x_{1},x_{2},...,x_{N}\}$, Test set $T$, Multi-modal model $M$, Complexity scorer $C$, Number of iterations $k$, Sample size $n$
0: Final accuracy on test set $T$
1: Initialize response log $R=\{\emptyset\mid x_{i}\in D\}$ {Empty list for each sample}
2:for $i=1$ to $k$ do
3:  for each $x\in D$ do
4:   $y=M(x)$ {Get model’s response to sample xx}
5:   Append $y$ to $R[x]$ {Store response for sample xx}
6:  end for
7:end for
8: Initialize disagreement scores $Disagreement=\{0\mid x_{i}\in D\}$
9:for each $x\in D$ do
10:  $UniqueResponses=\text{Remove duplicates from }R[x]$
11:  $k^{\prime}=\text{Length of }UniqueResponses$
12:  $Disagreement[x]=k^{\prime}/k$
13:end for
[…]
16:for each $x\in HardSamples\cup EasySamples$ do
17:  $Complexity[x]=C(x)$ {Get complexity score from scorer}
18:end for
19: Sort HardSamplesHardSamples by $Complexity[x]$ in ascending order
20:$HardLow=$ First $\frac{n}{4}$ samples from sorted HardSamplesHardSamples {Low complexity hard samples}
21:$HardHigh=$ Last $\frac{n}{4}$ samples from sorted HardSamplesHardSamples {High complexity hard samples}
22: Sort EasySamplesEasySamples by $Complexity[x]$ in ascending order
23:$EasyLow=$ First $\frac{n}{4}$ samples from sorted EasySamplesEasySamples {Low complexity easy samples}
24:$EasyHigh=$ Last $\frac{n}{4}$ samples from sorted EasySamplesEasySamples {High complexity easy samples}
[…]
26:$Correct=0$
27:for each $t\in T$ do
28:  $y_{pred}=M(t\mid Exemplars)$ {Model prediction with exemplars}
29:  if $y_{pred}==$ true label of $t$ then
30:   $Correct=Correct+1$
31:  end if
32:end for
[…]
34:return AccuracyAccuracy